OpenStudy (anonymous):

What is the decibel level of a normal conversation which has an intensity of 3.2 × 10-6 watts/m2?

OpenStudy (aaronq):

\(\sf dB=10*log(\dfrac{I}{I_o})=10*log(\dfrac{3.2∗10^{−6}W/m^2}{1.0∗10^{−12}W/m^2})\)